# thanks: these authors contributed equally to this workthanks: these authors contributed equally to this workthanks: these authors contributed equally to this work

### v3

@@ Abstract
For arbitrary $n$-qubit system, $\forall i$, there always exists at least one stabilizer generator $g_{j}$, such that the $g_{j}$ does not contain an $I$ on the $i$th qubit. Otherwise the rest (n1)(n-1) qubits have $n$ stabilizers, and the vector space stabilized by this stabilizer group is $2^{(n-1)-n}$ which is unacceptable [39]. We consider a subgroup of $\mathcal{G}$ defined as ${\mathcal{G}_{j}}=\langle g_{0},g_{1},...,g_{j-1},g_{j+1},...,g_{n-1}\rangle$ and its coset $g_{j}\mathcal{G}_{j}$} with $g_{j}\mathcal{G}_{j}\cup\mathcal{G}_{j}=\mathcal{G}$. For $\forall S_{m}\in\mathcal{G}_{j}$, there exists a unique stabilizer $S_{m^{\prime}}\in g_{j}\mathcal{G}_{j}$ which satisfies $S_{m^{\prime}}=g_{j}S_{m}$. Since $g_{j}$ does not contain an $I$ on the $i$th qubit, $S_{m}$ and $S_{l}$ do not contain two $I$s on the $i$th qubit at the same time. Then $\forall S_{m},S_{l}\in\mathcal{G}_{j}$, if $S_{m}\neq S_{l}$, $S_{m^{\prime}}$ is not equal to $S_{l^{\prime}}$. Thus $\forall i$, the fraction $f_{i,I}$, defined as the ratio of stabilizers containing $I$ on the $i$th qubit to the total number of stabilizers, satisfies $f_{i,I}\leq 1/2$. The state fidelity $\tilde{\mathcal{F}}$ is then estimated to be
@@ Appendix A An introduction to the conventional Qauntum Readout Error Mitigation (QREM)
The vector $(1,-1,-1,1)^{\rm T}$, $(1,1,-1,-1)^{\rm T}$, and $(1,-1,1,-1)^{\rm T}$ are unnormalized eigenvectors of $(\Lambda_{0}\otimes\Lambda_{1})(M_{0}\otimes M_{1})$, corresponding to the observables $ZZ$, ZIZI, and $IZ$, respectively. From the derivation above, it is cleared that the QREM process fixes the readout error perfectly and try to correct the initialization error by magnify the result with a factor of $(1-2q_{i})^{-1}$ for each measured qubit.
@@ Appendix B The derivation of entanglement state fidelity overestimation
In the QREM, if readout errors and preparation errors are not distinguished, the corrected result invariably forces $\lambda_{k}$ to $1$ while simultaneously neglects $\{q_{i}\}$. That is, if the measurement operator is $A_{k}$, the final result is tend to multiply the coefficient $1/\lambda_{k}$ [52].
@@ B.1 The linear graph state
The mapping $\{b_{k}\}$ is determined by $\{S_{k}\}$, Any operation other than the identity $I$ requires a measurement. Thus, a $Z$ operator is applied to qubit $Q_{i}$ in the operation $A_{b_{k}}$, such as $b_{1}=3$ with $A_{1}=ZII\cdots I$, $S_{1}=XZII\cdots I$, and $A_{b_{1}}=A_{3}=ZZII\cdots I$. In the contribution $\lambda_{k}/\lambda_{b_{k}}$, the numerator is given by the number of $Z$ operators in $A_{k}$, and the denominator by that in $A_{b_{k}}$, $\lambda_{k}/\lambda_{b_{k}}=(1-2q)^{-t}$. The exponent $t$ characterizes the contribution of stabilizer to the overestimated fidelity $\tilde{\mathcal{F}}$.
@@ B.3 The fully-connect graph state
according to Eq. (17). The latter type contributes the ideal $1$ to the expectation value, while the former contribute $(1-2q)^{n-t}/(1-2q)^{n}$ with (nt)(n-t) operator $Z$ in $A_{k}$. Therefore,
@@ B.4 The linear GHZ state
where $c_{i,k}=0$ if $g_{i}$ contains an $I$ on the $k$th qubit, otherwise $c_{i,k}=1$. Equation (32) imposes a very strong restriction, in other words, $\sum_{k}c_{i,k}=1$. If $\forall i,c_{i,k}=\delta_{ik}$, in other words, $g_{i}$ contains only one non-$I$ Pauli operator on the $i$th qubit and contains $I$s on all the other qubits, leading to the Clifford circuit with only single-qubit operations. Otherwise, the equation $q_{i}=q_{k}$ is necessary, while the Clifford circuit also consists of single-qubit operations equivalently with SWAP operations. No matter which situation it is, the system is thus a tensor product of single qubit states.
@@ Appendix D Implementation of Variational Quantum EigenSolver and Quantum Time Evolution for Chemistry Systems
where $\{\sigma^{\alpha}_{p},\sigma^{\beta}_{q},\cdots\}$ are Pauli operators $\{\sigma_{X},\sigma_{y},\sigma_{z},I\}$ on orbitals $\{\alpha,\beta,\cdots,p,q,\dots\}$, and $\{\tilde{\theta}\}$ and $\{\theta\}$ span the same parameter space. These unitary operators are thus decomposed using approximation schemes such as Trotter-Suzuki decomposition [19, 4] and mapped to quantum circuits, leading to a number of approximately $\mathcal{O}(N^{4})\sim\mathcal{O}(N^{5})$ gates where $N$ is the number of orbitals.
@@ 26      C += $\rm{HY}_{i}$
In the benchmark for quantum time evolution, a hydrogen molecule with bond length r(H-H)=2.0 Å and cc-pVTZ basis set is used under state preparation errors and Trotter approximation. Different active spaces are constructing using the lowest 2 to 16 orbitals, leading to quantum circuits with qubit counts ranging from 4 to 32. The evaluated energy for this systems is

### v4

@@ Abstract
For arbitrary $n$-qubit system, $\forall i$, there always exists at least one stabilizer generator $g_{j}$, such that the $g_{j}$ does not contain an II on the $i$th qubit. Otherwise the rest (n1)(n-1) qubits have $n$ stabilizers, and the vector space stabilized by this stabilizer group is $2^{(n-1)-n}$ which is unacceptable [39]. We consider a subgroup of $\mathcal{G}$ defined as ${\mathcal{G}_{j}}=\langle g_{0},g_{1},...,g_{j-1},g_{j+1},...,g_{n-1}\rangle$ and its coset $g_{j}\mathcal{G}_{j}$} with $g_{j}\mathcal{G}_{j}\cup\mathcal{G}_{j}=\mathcal{G}$. For $\forall S_{m}\in\mathcal{G}_{j}$, there exists a unique stabilizer $S_{m^{\prime}}\in g_{j}\mathcal{G}_{j}$ which satisfies $S_{m^{\prime}}=g_{j}S_{m}$. Since $g_{j}$ does not contain an II on the $i$th qubit, $S_{m}$ and $S_{m^{\prime}}$ do not contain two IIs on the $i$th qubit at the same time. Then $\forall S_{m},S_{l}\in\mathcal{G}_{j}$, if $S_{m}\neq S_{l}$, $S_{m^{\prime}}$ is not equal to $S_{l^{\prime}}$. Thus $\forall i$, the fraction $f_{i,I}$, defined as the ratio of stabilizers containing II on the $i$th qubit to the total number of stabilizers, satisfies $f_{i,I}\leq 1/2$. The state fidelity $\tilde{\mathcal{F}}$ is then estimated to be
@@ Appendix A An introduction to the conventional Qauntum Readout Error Mitigation (QREM)
The vector $(1,-1,-1,1)^{\rm T}$, $(1,1,-1,-1)^{\rm T}$, and $(1,-1,1,-1)^{\rm T}$ are unnormalized eigenvectors of $(\Lambda_{0}\otimes\Lambda_{1})(M_{0}\otimes M_{1})$, corresponding to the observables ZZZZ, ZIZI, and IZIZ, respectively. From the derivation above, it is cleared that the QREM process fixes the readout error perfectly and try to correct the initialization error by magnify the result with a factor of $(1-2q_{i})^{-1}$ for each measured qubit.
@@ Appendix B The derivation of entanglement state fidelity overestimation
In the QREM, if readout errors and preparation errors are not distinguished, the corrected result invariably forces $\lambda_{k}$ to 11 while simultaneously neglects $\{q_{i}\}$. That is, if the measurement operator is $A_{k}$, the final result is tend to multiply the coefficient $1/\lambda_{k}$ [52].
@@ B.1 The linear graph state
The mapping $\{b_{k}\}$ is determined by $\{S_{k}\}$, Any operation other than the identity II requires a measurement. Thus, a $Z$ operator is applied to qubit $Q_{i}$ in the operation $A_{b_{k}}$, such as $b_{1}=3$ with $A_{1}=ZII\cdots I$, $S_{1}=XZII\cdots I$, and $A_{b_{1}}=A_{3}=ZZII\cdots I$. In the contribution $\lambda_{k}/\lambda_{b_{k}}$, the numerator is given by the number of $Z$ operators in $A_{k}$, and the denominator by that in $A_{b_{k}}$, $\lambda_{k}/\lambda_{b_{k}}=(1-2q)^{-t}$. The exponent $t$ characterizes the contribution of stabilizer to the overestimated fidelity $\tilde{\mathcal{F}}$.
@@ B.3 The fully-connect graph state
according to Eq. (17). The latter type contributes the ideal 11 to the expectation value, while the former contribute $(1-2q)^{n-t}/(1-2q)^{n}$ with (nt)(n-t) operator $Z$ in $A_{k}$. Therefore,
@@ B.4 The linear GHZ state
where $c_{i,k}=0$ if $g_{i}$ contains an II on the $k$th qubit, otherwise $c_{i,k}=1$. Equation (32) imposes a very strong restriction, in other words, $\sum_{k}c_{i,k}=1$. If $\forall i,c_{i,k}=\delta_{ik}$, in other words, $g_{i}$ contains only one non-II Pauli operator on the $i$th qubit and contains IIs on all the other qubits, leading to the Clifford circuit with only single-qubit operations. Otherwise, the equation $q_{i}=q_{k}$ is necessary, while the Clifford circuit also consists of single-qubit operations equivalently with SWAP operations. No matter which situation it is, the system is thus a tensor product of single qubit states.
@@ Appendix D Implementation of Variational Quantum EigenSolver and Quantum Time Evolution for Chemistry Systems
where $\{\sigma^{\alpha}_{p},\sigma^{\beta}_{q},\cdots\}$ are Pauli operators $\{\sigma_{X},\sigma_{y},\sigma_{z},I\}$ on orbitals $\{\alpha,\beta,\cdots,p,q,\dots\}$, and $\{\tilde{\theta}\}$ and $\{\theta\}$ span the same parameter space. These unitary operators are thus decomposed using approximation schemes such as Trotter-Suzuki decomposition [19, 4] and mapped to quantum circuits, leading to a number of approximately $\mathcal{O}(N^{4})\sim\mathcal{O}(N^{5})$ gates where NN is the number of orbitals.
@@ 26      C += $\rm{HY}_{i}$
In the benchmark for quantum time evolution, a hydrogen molecule with bond length r(H-H)=2.0 Å and $c$-pVTZ basis set is used under state preparation errors and Trotter approximation. Different active spaces are constructing using the lowest 2 to 16 orbitals, leading to quantum circuits with qubit counts ranging from 4 to 32. The evaluated energy for this systems is